This is the first equation,x= floor plan 1, y=floor plan 2 with total of 38 houses,
x+y=38
now I need an equation to illustrate this, floor plan 1 sells for 175,000, and floor plan 2 sells for 200,000. the available houses combined are worth 7,200,000
then using the elimination method to determine how many houses with each floor plan are available
I would like to see how you do this step by step, and thank you very much.

Answer by rfer(16322)

x+y=38, x=38-y
-----------------
175000x+200000y=7200000
----------------------
175000(38-y)+200000y=7200000
6650000-175000y+200000y=7200000
25000y=550000
y=22 more expensive
x=16 cheaper house
----------------------------------
22*200000=4400000
16*175000=2800000